# An introduction to infinite projected entangled-pair state methods for variational ground state simulations using automatic differentiation

### v3

@@ III.1 Honeycomb lattice
We will now describe the general technical framework to simulate honeycomb lattices with the backbone CTMRG procedure described in Sec. II.2. To this end we consider an elementary unit cell of the honeycomb lattice. Here we choose to define it along so-called $x$-links for reasons that become clear soon. Alternatively and equivalently, it could as well be defined along $y$- or $z$-links. As an example with eight different tensors on the honeycomb lattice, corresponding to four elementary unit cells is shown in Fig. 18.
@@
III.2 Kagome lattice
Figure 20: Regular Kagome lattice with corner-sharing triangles and an elementary unit cell consisting of three lattice sites.
Another important and often encountered lattice in condensed matter physics is the Kagome lattice. It is of special interest due to its corner-sharing triangles, which lead a strong geometric frustration for anti-ferromagnetic models. Using a simple mapping of the Kagome lattice to a square lattice, we can directly incorporate it into our variational PEPS library. The Kagome lattice is shown in Fig. 20. Naturally, we can define a unit cell of tensors that is repeated periodically over the whole two-dimensional lattice. In our setting we consider an upward triangle on the Kagome lattice as an elementary unit cell, highlighted by the gray dotted area in Fig. 20. By choosing a coarse-graining, we can represent the three lattice sites in the unit cell by a single iPEPS tensor, which connects to its neighbours by four virtual indices. This direct mapping is shown in Fig. 21.
Figure 21: Regular Kagome lattice mapped to a square lattice by coarse-graining of the three spins in each unit cell.
Nearest-neighbour links in the Kagome lattice get mapped to nearest-neighbour or second-nearest-neighbour links in the square lattice. Every iPEPS site on the square lattice has a physical dimension of $p^{3}$. As an alternative mapping, which results in the same coarse-grained TN structure, we move from the Kagome lattice to its dual, the honeycomb lattice. Here the spins live on the links instead of the vertices. The honeycomb mapping presented in Sec. III.1 is therefore not directly applicable and additional simplex tensors are necessary to connect the lattices sites. This TN structure is shown in Fig. 22, which is commonly known as the infinite projected entangled simplex state (iPESS) [95].
@@ III.2 Kagome lattice
Due to this particular mapping, three Kagome lattice sites (along with two simplex tensors) are coarse-grained into a single iPEPS site on the square lattice. While the mappings in Fig. 21 and Fig. 22 result in the same square lattice TN, they differ in the number of variational parameters in the ansatz. In the direct iPEPS ansatz, every unit cell tensor has $p^{3}\chi_{B}^{4}$ parameters, while there are only $(3p\chi_{B}^{2}+2\chi_{B}^{3})$ parameters for the iPESS ansatz. Moreover, quantum correlations between lattice sites are exactly captured within the coarse-grained cluster for the iPEPS, whereas they are limited by the bulk bond dimension for the iPESS. In the ladder case, however, there is no bias between lattice sites within one cluster and sites belonging to different clusters. The nearest neighbor interactions on the Kagome lattice are mapped to on-site, nearest neighbor and next-nearest neighbor interactions on the square lattice. As a concrete mapping example which has particular use in the study of the regular Heisenberg model in a magnetic field, we consider the iPEPS configuration
@@
III.3 Square-Kagome lattice
@@ III.4 Triangular lattice
III.5 Comments about different structures
@@ IV Benchmarks and discussions
where $\langle i,j\rangle$ denotes nearest neighbours and $\vec{S}_{i}$ are the spin-$1/2$ operators on the lattice sites. We consider isotropic anti-ferromagnetic interactions at $J=1.0$ throughout the benchmark section. Variational energies obtained with our implementation are denoted by “variational update” (VU). Where applicable, we include different TN variants (e.g., iPESS and iPEPS) in the numerical benchmarks, to highlight the effect of different numbers of variational parameters. Imaginary time-evolution in the form of a “simple update” (SU) on the different lattice structures can provide initial states for the variational update as discussed in Sec. II.7.3. Whenever we use initial tensors from the SU, we add a small amount of random noise to the input tensors prior to the variational update, in order to circumvent possible local minima in the imaginary time evolution.
@@
We add for each benchmark a table with the comparison of the results obtained by the simple update simulations and the best result throughout all variational updates for a fixed iPEPS bond dimension $\chi_{B}$. Both expecation values have been calculated by CTMRG.
IV.1 Comments on lower bounds in variational principles
@@ IV.3 Kagome lattice
Since the ground state is known to be a spin liquid state, that does not form any magnetic ordering down to zero temperature while preserving lattice translation and rotation symmetry, we use the smallest unit cells of only three sites in our simulations. The SU then works on the three-site iPESS ansatz. The VU is performed both on the honeycomb iPESS and on a coarse-grained, fully translationally invariant iPEPS state. The number of variational parameters are hence $(3p\chi_{B}^{2}+2\chi_{B}^{3})$ for the iPESS and $p^{3}\chi_{B}^{4}$ for the iPEPS. Again, the iPEPS state is more expressive and produces lower variational energies, that follow a smoother convergence with bond dimension $\chi_{B}$, see Fig. 31. The ED energy provides a lower-bound for the energy, as argued in Sec. IV.1. Our energies are compatible with other state-of-the-art numerical methods as the extrapolated iPESS result from Ref. [103], but we would like to point out that the authors noted that their results are not variational and hence the comparison is slightly tainted. Our result showcases the purpose of variational iPEPS optimization for highly frustrated systems to obtain a real upper bound to the ground state energy.
IV.4 Square-Kagome lattice
@@
Figure 32: Benchmarking results for the isotropic spin-$1/2$ Heisenberg model on the square-Kagome lattice. For comparison, we include the variational Monte-Carlo results presented in Ref. [108]. Additionally, we show the extrapolated iPEPS result obtained in Ref. [109], which, to be strict, is not variational. We stress that the mean-field energies also are not variational as discussed in Sec. IV.
@@ IV.5 Triangular lattice
IV.6 Comments on excited states
@@
IV.7 Comments on fermionic systems
@@ V Conclusion and prospects
V.1 $\text{CO}_{2}$-emissions table
@@
Acknowledgements.
We acknowledge inspiring discussions with Ji-Yao Chen, Andreas Haller, Juraj Hasik, Augustine Kshetrimayum, Alexander Nietner and Niklas Tausendpfund. We would like to particularly thank Boris Ponsioen, who has for shared valuable insights, and Frederik Wilde, who has helped us to get the details of the automatic differentiation and the custom fixed-point derivative correct. E. L. W. thanks the Studienstiftung des deutschen Volkes for support. This work has been funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the project number 277101999 – CRC 183 (project B01), for which this constitutes an inter-node publication involving both Cologne and Berlin, and the BMBF (MUNIQC-Atoms, FermiQP). It has also received funding from the Cluster of Excellence MATH+ and from the Quantum Flagship (PasQuans2). We would like to thank the ZEDV (IT support) of the physics department, Freie Universität Berlin, for computing time and their technical support, particularly we thank Jörg Behrmann and Jens Dreger. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at the Jülich Supercomputing Centre (JSC) (Grant NeTeNeSyQuMa) and the FZ Jülich for JURECA (institute project PGI-8) [127]. For the Python version we thank the developers of the JAX framework [128, 129] for their work to provide an AD framework with optimized numerical operations and their technical support during the development of this work. We also acknowledge the use of the TensorKit package [130] in the Julia version of the code and wish to advertise the open source libraries of the Quantum Ghent group in this context [131]. We make use of the Zygote [132] package for AD in the Julia programming language. The work has been discussed and refined during the workshop “Tensor Networks: Mathematical Structures and Novel Algorithms (2022)” at the Erwin Schrödinger International Institute for Mathematics and Physics in Vienna and the workshop “Entanglement in Strongly Correlated Systems (2023)” at the Centro de Ciencias de Benasque Pedro Pascual. We thank the organizers for their hospitality and work.
@@ Appendix A Adjoint functions and variables
$\begin{split}h&=h_{2}\circ h_{1},\\ h_{1}&:M_{n\times n}\times M_{n\times n}\xrightarrow{}M_{n\times n},\\ h_{2}&:M_{n\times n}\xrightarrow{}\mathbb{R},\end{split}$ (27)
@@
$dx=\frac{\partial h_{2}}{\partial C}dC=:\sum_{i,j}\bar{C}_{i,j}dC_{i,j}=\mathrm{Tr}(\bar{C}^{\mkern-1.5mu\mathsf{T}}dC).$ (28)
In the first equation, we have suppressed the sum over the indices of $C$. Eq. (28) defines the adjoint variable $\bar{C}$ of $C$. We see that the adjoint variable $\bar{C}$ is the derivative of the scalar output of the function $h_{2}$ w.r.t. $C$. Thus, for the case of a scalar output the variable $C$ and the adjoint variable $\bar{C}$ have the same dimension. Now, in order to get the gradient $\nabla h$ we are interested in the derivative of the output w.r.t. the input variables $(A,B)$. To this end we consider the differential of the intermediate variable
$dC=\frac{\partial h_{1}}{\partial A}dA+\frac{\partial h_{1}}{\partial B}dB.$ (29)
Inserting this into Eq. (28), we obtain
$dx=\mathrm{Tr}\bigg{(}\underbrace{\bar{C}^{\mkern-1.5mu\mathsf{T}}\frac{\partial h_{1}}{\partial A}}_{\bar{A}^{\mkern-1.5mu\mathsf{T}}}dA\bigg{)}+\mathrm{Tr}\bigg{(}\underbrace{\bar{C}^{\mkern-1.5mu\mathsf{T}}\frac{\partial h_{1}}{\partial B}}_{\bar{B}^{\mkern-1.5mu\mathsf{T}}}dB\bigg{)}.$ (30)
@@
$\bar{h}_{1}:\bar{C}^{\mkern-1.5mu\mathsf{T}}\xmapsto{}(\bar{A}^{\mkern-1.5mu\mathsf{T}},\bar{B}^{\mkern-1.5mu\mathsf{T}})^{\mkern-1.5mu\mathsf{T}}=\left(\bar{C}^{\mkern-1.5mu\mathsf{T}}\frac{\partial h_{1}}{\partial A},\bar{C}^{\mkern-1.5mu\mathsf{T}}\frac{\partial h_{1}}{\partial B}\right)^{\mkern-1.5mu\mathsf{T}}.$ (31)
@@
$\begin{split}\nabla h=\left(\bar{A},\bar{B}\right)&=\left(\frac{\partial h_{1}}{\partial A}\bar{C},\frac{\partial h_{1}}{\partial B}\bar{C}\right)\\ &=\left(\frac{\partial h_{1}}{\partial A}\frac{\partial h_{2}}{\partial C},\frac{\partial h_{1}}{\partial B}\frac{\partial h_{2}}{\partial C}\right).\end{split}$ (32)
@@ Appendix C The implicit function theorem and its use at the CTMRG fixed-point
$\partial_{A}e^{*}(A)=\partial_{A}c(A,e^{*})+\partial_{e^{*}}c(A,e^{*})\partial_{A}e^{*}(A).$ (33)
@@
$\begin{split}\partial_{A}e^{*}(A)=&(L+K\partial_{A}e^{*}(A))\\ =&\left(\sum_{n=0}^{\infty}K\right)L=(\mathds{1}-K)^{-1}L.\end{split}$ (34)
As discussed in Appendix A, we aim at finding the adjoint function of the CTMRG iteration at the fixed point, which is a vector-Jacobian product (vJP) $\mathbf{v}^{\mkern-1.5mu\mathsf{T}}\partial_{A}e^{*}(A)$. Inserting Eq. (34) yields
$\mathbf{v}^{\mkern-1.5mu\mathsf{T}}\partial e^{*}(A)=\mathbf{v}^{\mkern-1.5mu\mathsf{T}}(\mathds{1}-K)^{-1}L=\mathbf{w}^{\mkern-1.5mu\mathsf{T}}L,$ (35)
@@
$\mathbf{w}^{\mkern-1.5mu\mathsf{T}}=\mathbf{v}^{\mkern-1.5mu\mathsf{T}}+\mathbf{w}^{\mkern-1.5mu\mathsf{T}}K.$ (36)
Here $\mathbf{w}^{\mkern-1.5mu\mathsf{T}}K$ is another vJP but this time only dependent of the derivative of a single absorption step evaluated at the fixed-point of the CTMRG routine. Solving Eq. (36) we can find $\mathbf{w}^{\mkern-1.5mu\mathsf{T}}$ to calculate the vJP of the CTMRG routine from Eq. (35). In the end we reduced the naive effort of unrolling the fixed-point iterations to just calculate the derivative of a single CTMRG iteration and another fixed-point iteration which both are much less memory intensive.

### v4

@@ 3.1 Honeycomb lattice
We will now describe the general technical framework to simulate honeycomb lattices with the backbone CTMRG procedure described in Sec. 2.2. To this end we consider an elementary unit cell of the honeycomb lattice. Here we choose to define it along so-called $x$-links for reasons that become clear soon. Alternatively and equivalently, it could as well be defined along $y$- or $z$-links. As an example with eight different tensors on the honeycomb lattice, corresponding to four elementary unit cells is shown in Fig. 18.
@@
3.2 Kagome lattice
Figure 20: (a) Regular Kagome lattice with corner-sharing triangles and an elementary unit cell consisting of three lattice sites. (b) Regular Kagome lattice mapped to a square lattice by coarse-graining of the three spins in each unit cell.
Another important and often encountered lattice in condensed matter physics is the Kagome lattice. It is of special interest due to its corner-sharing triangles, which lead a strong geometric frustration for anti-ferromagnetic models. Using a simple mapping of the Kagome lattice to a square lattice, we can directly incorporate it into our variational PEPS library. The Kagome lattice is shown in Fig. 20(a). Naturally, we can define a unit cell of tensors that is repeated periodically over the whole two-dimensional lattice. In our setting we consider an upward triangle on the Kagome lattice as an elementary unit cell, highlighted by the gray dotted area in Fig. 20(a). By choosing a coarse-graining, we can represent the three lattice sites in the unit cell by a single iPEPS tensor, which connects to its neighbours by four virtual indices. This direct mapping is shown in Fig. 20(b). Nearest-neighbour links in the Kagome lattice get mapped to nearest-neighbour or second-nearest-neighbour links in the square lattice. Every iPEPS site on the square lattice has a physical dimension of $p^{3}$. As an alternative mapping, which results in the same coarse-grained TN structure, we move from the Kagome lattice to its dual, the honeycomb lattice. Here the spins live on the links instead of the vertices. The honeycomb mapping presented in Sec. 3.1 is therefore not directly applicable and additional simplex tensors are necessary to connect the lattices sites. This TN structure is shown in Fig. 21, which is commonly known as the infinite projected entangled simplex state (iPESS) [108].
@@ 3.2 Kagome lattice
Due to this particular mapping, three Kagome lattice sites (along with two simplex tensors) are coarse-grained into a single iPEPS site on the square lattice. While the mappings in Fig. 20(b) and Fig. 21 result in the same square lattice TN, they differ in the number of variational parameters in the ansatz. In the direct iPEPS ansatz, every unit cell tensor has $p^{3}\chi_{B}^{4}$ parameters, while there are only $(3p\chi_{B}^{2}+2\chi_{B}^{3})$ parameters for the iPESS ansatz. Moreover, quantum correlations between lattice sites are exactly captured within the coarse-grained cluster for the iPEPS, whereas they are limited by the bulk bond dimension for the iPESS. In the ladder case, however, there is no bias between lattice sites within one cluster and sites belonging to different clusters. The nearest neighbor interactions on the Kagome lattice are mapped to on-site, nearest neighbor and next-nearest neighbor interactions on the square lattice. As a concrete mapping example which has particular use in the study of the regular Heisenberg model in a magnetic field, we consider the iPEPS configuration
@@
3.3 Square-Kagome lattice
@@ 3.4 Triangular lattice
3.5 Comments about different structures
@@ 4 Benchmarks and discussions
where $\langle i,j\rangle$ denotes nearest neighbours and $\vec{S}_{i}$ are the spin-$1/2$ operators on the lattice sites. We consider isotropic anti-ferromagnetic interactions at $J=1.0$ throughout the benchmark section. Variational energies obtained with our implementation are denoted by “variational update” (VU). Where applicable, we include different TN variants (e.g., iPESS and iPEPS) in the numerical benchmarks, to highlight the effect of different numbers of variational parameters. Imaginary time-evolution in the form of a “simple update” (SU) on the different lattice structures can provide initial states for the variational update as discussed in Sec. 2.8.3. Whenever we use initial tensors from the SU, we add a small amount of random noise to the input tensors prior to the variational update, in order to circumvent possible local minima in the imaginary time evolution.
@@
We add for each benchmark a table with the comparison of the results obtained by the simple update simulations and the best result throughout all variational updates for a fixed iPEPS bond dimension $\chi_{B}$. Both expectation values have been calculated by CTMRG.
4.1 Comments on lower bounds in variational principles
@@ 4.3 Kagome lattice
Since the ground state is known to be a spin liquid state, that does not form any magnetic ordering down to zero temperature while preserving lattice translation and rotation symmetry, we use the smallest unit cells of only three sites in our simulations. The SU then works on the three-site iPESS ansatz. The VU is performed both on the honeycomb iPESS and on a coarse-grained, fully translationally invariant iPEPS state. The number of variational parameters are hence $(3p\chi_{B}^{2}+2\chi_{B}^{3})$ for the iPESS and $p^{3}\chi_{B}^{4}$ for the iPEPS. Again, the iPEPS state is more expressive and produces lower variational energies, that follow a smoother convergence with bond dimension $\chi_{B}$, see Fig. 30. The ED energy provides a lower-bound for the energy, as argued in Sec. 4.1. Our energies are compatible with other state-of-the-art numerical methods as the extrapolated iPESS result from Ref. [116], but we would like to point out that the authors noted that their results are not variational and hence the comparison is slightly tainted. Our result showcases the purpose of variational iPEPS optimization for highly frustrated systems to obtain a real upper bound to the ground state energy.
4.4 Square-Kagome lattice
@@
Figure 31: Benchmarking results for the isotropic spin-$1/2$ Heisenberg model on the square-Kagome lattice. For comparison, we include the variational Monte-Carlo results presented in Ref. [121]. Additionally, we show the extrapolated iPEPS result obtained in Ref. [122], which, to be strict, is not variational. We stress that the mean-field energies also are not variational as discussed in Sec. 4.
@@ 4.5 Triangular lattice
4.6 Comments on excited states
@@
4.7 Comments on fermionic systems
@@ 5 Conclusion and prospects
5.1 $\text{CO}_{2}$-emissions table
@@
Acknowledgements
We acknowledge inspiring discussions with Ji-Yao Chen, Andreas Haller, Juraj Hasik, Augustine Kshetrimayum, Alexander Nietner and Niklas Tausendpfund. We would like to particularly thank Boris Ponsioen, who has for shared valuable insights, and Frederik Wilde, who has helped us to get the details of the automatic differentiation and the custom fixed-point derivative correct. We would like to thank the ZEDV (IT support) of the physics department, Freie Universität Berlin, for computing time and their technical support, particularly we thank Jörg Behrmann and Jens Dreger.
For the Python version we thank the developers of the JAX framework [140, 141] for their work to provide an AD framework with optimized numerical operations and their technical support during the development of this work. We also acknowledge the use of the TensorKit package [91] in the Julia version of the code and wish to advertise the open source libraries of the Quantum Ghent group in this context [142]. We make use of the Zygote [143] package for AD in the Julia programming language.
The work has been discussed and refined during the workshop “Tensor Networks: Mathematical Structures and Novel Algorithms (2022)” at the Erwin Schrödinger International Institute for Mathematics and Physics in Vienna and the workshop “Entanglement in Strongly Correlated Systems (2023)” at the Centro de Ciencias de Benasque Pedro Pascual. We thank the organizers for their hospitality and work.
Funding information
E. L. W. thanks the Studienstiftung des deutschen Volkes for support. This work has been funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the project number 277101999 – CRC 183 (project B01), for which this constitutes an inter-node publication involving both Cologne and Berlin, and the BMBF (MUNIQC-Atoms, FermiQP). It has also received funding from the Cluster of Excellence MATH+ and from the Quantum Flagship (PasQuans2).
The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at the Jülich Supercomputing Centre (JSC) (Grant NeTeNeSyQuMa) and the FZ Jülich for JURECA (institute project PGI-8) [144].
@@ Appendix A Adjoint functions and variables
$\begin{split}h&=h_{2}\circ h_{1},\\ h_{1}&:M_{n\times n}\times M_{n\times n}\xrightarrow{}M_{n\times n},\\ h_{2}&:M_{n\times n}\xrightarrow{}\mathbb{R},\end{split}$ (A.1)
@@
$dx=\frac{\partial h_{2}}{\partial C}dC=:\sum_{i,j}\bar{C}_{i,j}dC_{i,j}=\mathrm{Tr}(\bar{C}^{\mkern-1.5mu\mathsf{T}}dC).$ (A.2)
In the first equation, we have suppressed the sum over the indices of $C$. Eq. (A.2) defines the adjoint variable $\bar{C}$ of $C$. We see that the adjoint variable $\bar{C}$ is the derivative of the scalar output of the function $h_{2}$ w.r.t. $C$. Thus, for the case of a scalar output the variable $C$ and the adjoint variable $\bar{C}$ have the same dimension. Now, in order to get the gradient $\nabla h$ we are interested in the derivative of the output w.r.t. the input variables $(A,B)$. To this end we consider the differential of the intermediate variable
$dC=\frac{\partial h_{1}}{\partial A}dA+\frac{\partial h_{1}}{\partial B}dB.$ (A.3)
Inserting this into Eq. (A.2), we obtain
$dx=\mathrm{Tr}\bigg{(}\underbrace{\bar{C}^{\mkern-1.5mu\mathsf{T}}\frac{\partial h_{1}}{\partial A}}_{\bar{A}^{\mkern-1.5mu\mathsf{T}}}dA\bigg{)}+\mathrm{Tr}\bigg{(}\underbrace{\bar{C}^{\mkern-1.5mu\mathsf{T}}\frac{\partial h_{1}}{\partial B}}_{\bar{B}^{\mkern-1.5mu\mathsf{T}}}dB\bigg{)}.$ (A.4)
@@
$\bar{h}_{1}:\bar{C}^{\mkern-1.5mu\mathsf{T}}\xmapsto{}(\bar{A}^{\mkern-1.5mu\mathsf{T}},\bar{B}^{\mkern-1.5mu\mathsf{T}})^{\mkern-1.5mu\mathsf{T}}=\left(\bar{C}^{\mkern-1.5mu\mathsf{T}}\frac{\partial h_{1}}{\partial A},\bar{C}^{\mkern-1.5mu\mathsf{T}}\frac{\partial h_{1}}{\partial B}\right)^{\mkern-1.5mu\mathsf{T}}.$ (A.5)
@@
$\begin{split}\nabla h=\left(\bar{A},\bar{B}\right)&=\left(\frac{\partial h_{1}}{\partial A}\bar{C},\frac{\partial h_{1}}{\partial B}\bar{C}\right)\\ &=\left(\frac{\partial h_{1}}{\partial A}\frac{\partial h_{2}}{\partial C},\frac{\partial h_{1}}{\partial B}\frac{\partial h_{2}}{\partial C}\right).\end{split}$ (A.6)
@@ Appendix C The implicit function theorem and its use at the CTMRG fixed-point
$\partial_{A}e^{*}(A)=\partial_{A}c(A,e^{*})+\partial_{e^{*}}c(A,e^{*})\partial_{A}e^{*}(A).$ (C.1)
@@
$\begin{split}\partial_{A}e^{*}(A)=&(L+K\partial_{A}e^{*}(A))\\ =&\left(\sum_{n=0}^{\infty}K\right)L=(\mathds{1}-K)^{-1}L.\end{split}$ (C.2)
As discussed in Appendix A, we aim at finding the adjoint function of the CTMRG iteration at the fixed point, which is a vector-Jacobian product (vJP) $\mathbf{v}^{\mkern-1.5mu\mathsf{T}}\partial_{A}e^{*}(A)$. Inserting Eq. (C.2) yields
$\mathbf{v}^{\mkern-1.5mu\mathsf{T}}\partial e^{*}(A)=\mathbf{v}^{\mkern-1.5mu\mathsf{T}}(\mathds{1}-K)^{-1}L=\mathbf{w}^{\mkern-1.5mu\mathsf{T}}L,$ (C.3)
@@
$\mathbf{w}^{\mkern-1.5mu\mathsf{T}}=\mathbf{v}^{\mkern-1.5mu\mathsf{T}}+\mathbf{w}^{\mkern-1.5mu\mathsf{T}}K.$ (C.4)
Here $\mathbf{w}^{\mkern-1.5mu\mathsf{T}}K$ is another vJP but this time only dependent of the derivative of a single absorption step evaluated at the fixed-point of the CTMRG routine. Solving Eq. (C.4) we can find $\mathbf{w}^{\mkern-1.5mu\mathsf{T}}$ to calculate the vJP of the CTMRG routine from Eq. (C.3). In the end we reduced the naive effort of unrolling the fixed-point iterations to just calculate the derivative of a single CTMRG iteration and another fixed-point iteration which both are much less memory intensive.